# CREA: A Collaborative Multi-Agent Framework for Creative Content Generation with Diffusion Models

### v1

@@ Abstract
Creativity in AI imagery remains a fundamental challenge, requiring not only the generation of visually compelling content but also the capacity to add novel, expressive, and artistically rich transformations to images. Unlike conventional editing tasks that rely on direct prompt-based modifications, creative image editing demands an autonomous, iterative approach that balances originality, coherence, and artistic intent. To address this, we introduce CREA, a novel multi-agent collaborative framework that mimics the human creative process. Our framework leverages a team of specialized AI agents who dynamically collaborate to conceptualize, generate, critique, and enhance images. Through extensive qualitative and quantitative evaluations, we demonstrate that CREA significantly outperforms state-of-the-art methods in diversity, semantic alignment, and creative transformation. By structuring creativity as a dynamic, agentic process, CREA redefines the intersection of AI and art, paving the way for autonomous AI-driven artistic exploration, generative design, and human-AI co-creation. To the best of our knowledge, this is the first work to introduce the task of creative editing.
@@ 1 Introduction
Generative AI has significantly transformed the field of image generation, producing high-quality visuals with remarkable detail and realism. Advances in diffusion models [57, 14], GANs [31], and retrieval-augmented techniques [4, 65] have enabled powerful capabilities in content synthesis, making AI-driven image creation an essential tool for artists, designers, and various creative industries. These models have been widely applied in diverse tasks, including image-to-image translation [29, 52, 63], inpainting [57, 42, 30], style transfer [32, 20, 58, 79], and content-aware editing [46, 7, 6, 11], and opening up new frontiers in digital art, advertising, and entertainment, revolutionizing creative workflows.
@@
To address this limitation, we propose the task of creative image editing, where the goal is to modify images in a way that enhances their creative and artistic qualities with minimal user intervention. Unlike conventional editing tasks that focus on making explicit, text-driven modifications, creative editing aims to transform images into novel, aesthetically rich compositions in a disentangled way. Our approach mimics the complex process of creative image generation and editing as a collaborative team effort, drawing inspiration from human workflows where specialists iteratively refine ideas to achieve a shared artistic vision. To achieve this, we introduce a novel collaborative multi-agent framework, where agents, each with a distinct role such as Creative Director or Art Critic work in synergy to conceptualize, generate and refine creative outputs by grounding them in well-established creativity principles distilled from research from distinguished creativity researchers as shown in Table 4. This structured yet flexible approach enables the disentangled creation of highly diverse and imaginative images, ensuring both novelty and coherence at every stage of the generation process. Since creativity is a complicated task to assess and evaluate, our agents leverage the extensive knowledge in multimodal LLMs [51] based on state-of-the-art creativity principles inspired by previous work [5, 23, 54, 16, 1, 3]. Through both qualitative and quantitative evaluations, we demonstrate that our method consistently produces edits perceived as more creative and aesthetically pleasing compared to baseline methods. By dynamically infusing generative processes with artistic influences, our method expands the boundaries of creative exploration in image editing, offering new avenues for generative artistry. Our contributions are as follows:
@@
We demonstrate the versatility of our method across image editing and generation, as well as its potential for enhancing personalization workflows.
We further highlight the adaptability of our approach by showcasing its potential applications in creative video generation.
@@ 2 Related Work
Text-to-Image Models. Recent advances in diffusion models [26, 57, 59] have revolutionized text-to-image synthesis, enabling high-fidelity image generation guided by textual prompts. Models such as DALLE-3 [50], SDXL [53], and Flux [14] demonstrate the ability to generate visually compelling images based on textual prompts [71, 69, 39]. While these models produce high-quality outputs, they lack a structured mechanism to enforce creativity principles in generation. Personalized generation approaches such as DreamBooth [58] and Textual Inversion [20] focus on fine-tuning text-to-image models for specific subjects or styles. However, these methods optimize for style and subject consistency rather than generating creative images.
@@
Creative Image Generation and Editing. Research on creative generation in AI has advanced through GANs [22, 32] and diffusion models, leveraging contrastive loss and diversity-based objectives to encourage novel synthesis beyond category constraints [13, 25, 60]. Recent works such as ProCreate [40] use energy-based repulsion to steer diffusion models away from reference images, while Inspiration Tree [66] employs hierarchical decomposition for conceptual hybridization. ConceptLab [55] tackles creative concept synthesis through diffusion priors, iteratively enforcing constraints to generate novel category members. Despite their advancements, these methods require expensive retraining or optimization, and fail to generalize to broader concepts. For image editing, ControlNet [77] extends Stable Diffusion, introducing external conditioning signals for localized modifications while maintaining structural integrity. SDEdit [46] and Blended Latent Diffusion [2] refine diffusion-based editing for finer control, while InstructPix2Pix [7] allows text-based transformations via user prompts. However, these approaches focus on structural fidelity rather than creativity-driven, minimal-guidance transformations.
@@
Large Language Models Large Language Models (LLMs) such as GPT-4 [51] and PaLM [9] have demonstrated remarkable capabilities in natural language understanding and generative tasks. These models leverage transformer architectures [64], which allows them to process large amounts of text and generate coherent and contextually relevant outputs. In creative domains, LLMs have been increasingly integrated into AI-driven artistic workflows, assisting with idea generation, structured prompt synthesis, and artistic guidance [45]. GenArtist [68] uses a multimodal single agent for general image generation and editing but does not address creative image editing and fails to leverage the collaborative nature of multiple agents for complex use cases. Several studies have explored LLMs in the context of creativity [41, 80, 35, 10, 34, 18]; however, to the best of our knowledge, creative editing tasks have not been explicitly tackled before. Additionally, agentic frameworks—where AI systems exhibit autonomy in decision-making, iterative refinement, and adaptive goal-setting—have not yet been systematically explored in the context of creative image generation or editing.
@@ 3 Methodology
Creativity is a multifaceted process that involves ideation, refinement, evaluation, and the ability to balance novelty with coherence. Traditional generative AI systems, however, often rely on static, single-step processes, where outputs are determined by a fixed prompt-to-output pipeline, limiting adaptability and fine-grained control. In contrast, an agentic framework provides a more dynamic and interactive approach by coordinating multiple AI agents that can reason, collaborate, and iteratively refine creative outputs. By structuring creativity as a multi-agent process, our system mirrors human collaborative workflows, where distinct roles—such as conceptualization, critique, and enhancement—work in tandem to push creative boundaries while ensuring coherence and quality. We describe the key components of our framework below. Also, please see Fig. 2 and Algorithm 1 for an overview of our framework.
@@ 3.1 Multi-Agent Framework Design
Prompt Architect . Translates conceptual ideas into contrastive prompts [8] for each creativity principle, merging them into a high-creativity prompt for the Generative Executor agent for image editing/generation and refining them iteratively based on agent feedback.
Art Critic . The Art Critic evaluates the generated image based on the creativity principles, assigning a score for each criterion. Given that LLMs can approximate human judgment in subjective evaluations [81], this agent uses a multimodal LLM judge to ensure sensible evaluation.
@@
Each agent maintains a private memory and utilizes role-specific tools to share task status and relevant information, ensuring informed decision-making at every step. See Section C.4 in the Supplementary Material for detailed agent compositions.
@@ 3.2 Collaborative Agentic Synthesis
Creativity Principles. Our framework leverages six creativity principles, grounded in state-of-the-art creativity theories, to systematically evaluate and measure creative output. Originality, measuring novelty and uniqueness, is inspired by Boden’s Theory of Creativity [5] and Guilford’s Divergent Thinking [23] framework. Expressiveness, which captures emotional impact, is influenced by Amabile’s Model of Creativity [1] and Ramachandran & Hirstein’s Laws of Aesthetics [54]. Aesthetic Appeal, assessing composition and harmony, is grounded in Martindale’s Aesthetic Model [44] and Berlyne’s Aesthetic Theory [3]. Technical Execution, evaluating craftsmanship and skill, draws from Amabile’s Model and AI Creativity Frameworks [1]. Unexpected Associations, reflecting surprise and ingenuity, is supported by the Geneplore Model [17] and Boden’s Combinational Creativity [5]. Finally, Interpretability & Depth, which considers exploration potential, is informed by Ramachandran’s Laws [54] and the Geneplore Model [17] (see Appendix for more details). These six principles are used as a creativity template for agents to assess and refine the outputs.
3.2.1 Pre-Generation Planning and Image Synthesis
@@
The pre-generation planning phase serves as a structured ideation stage where agents collaboratively establish a creative blueprint, before image generation begins. This phase involves three key agents: the Creative Director , Prompt Architect , and Generative Executor , who collaboratively devise the creativity prompt that is both creatively rich and technically viable. First, Creative Director, interprets the initial image , either user-provided or generated using the user-provided concept to formulate a creativity blueprint , capturing the core theme, stylistic interpretation, visual structure, and necessary constraints to balance artistic flexibility with semantic coherence (see Appendix for more details). Based on , the Prompt Architect synthesizes a set of contrastive prompts , each conditioned on a distinct creativity principle and are merged into a high-creativity fused prompt through Chain-of-Thought [70] reasoning:
where CoT-Fusion extracts salient conceptual and stylistic attributes from each prompt and synthesizes them into a coherent, balanced, and conceptually rich formulation. The Generative Executor evaluates the feasibility of and determines T2I model-specific constraints-such as ControlNet conditioning scale, image guidance scale to anticipate the nuances of the generated blueprint. Once all agents reach a consensus, the prompt moves to the next phase for creative image generation.
@@
Image Generation: The Generative Executor, plays a dual role, either generating a creative image using a T2I model or performing disentangled creative editing on an existing image, whether initially generated () or user-provided (). The high-creativity structured prompt formulated during the pre-generation planning phase is then taken by the Generative Executor to synthesize the initial creative image, using a text-to-image diffusion model, as:
where represents model-specific parameters-such as guidance scale and ControlNet conditioning scale. The generated image, serves as the starting point for the creative editing and iterative refinement process as described in the following sections.
Image Editing: If the user provides an image instead of generating one, or if an enhancement to is required, the Generative Executor performs disentangled creative edits using
@@
where is the ControlNet model used to perform disentangled edits, is the high-creativity editing prompt generated in the pre-generation planning phase and represents parameters of . After generation or editing, the resulting image progresses to the post-generation evaluation phase to investigate if further iterative refinements are necessary.
3.2.2 Post-Generation Evaluation
@@
The post-generation evaluation phase assesses the edited creative image, (or for generation) against the creativity template, to maximize creativity. This Critic and the Creative Director collaborate, with the Critic systematically evaluating the initial edited image, based on the creativity template, . The Critic utilizes the LLM-as-a-Judge framework [81] to evaluate the edited image by assigning a creativity score, for each of the six creativity criteria on a 1-5 scale. The overall Creativity Index, is then computed as:
@@
If the total creativity score, , where is a predefined threshold, the edited image is considered suboptimal in creativity. The Creative Director then reviews the Critic’s evaluation and may challenge its assessment if the assigned scores misalign with the creative blueprint, . Once a consensus is reached, is either finalized or progresses to the next phase for further enhancement based on the final creativity score .
@@ 3.2.3 Self-Enhancement with Optional User-Guidance
While one round of editing produces optimal creative outputs, further refinement can enhance creativity. In the self-enhancement phase, all agents iteratively refine the edited image, with optional human intervention. Given a maximum of iterations, each edited image is evaluated by the Critic to compute its . When , the Refinement Stategist, identifies the low scoring creative dimensions and proposes a refinement plan to enhance the corresponding weak dimensions, which the Prompt Architect uses to refine the editing prompt, as follows:
where represents the prompt adjustment that amplify specific creativity dimensions based on their evaluation scores and is the refined creativity prompt. The refined prompt is used by the Generative Executor to regenerate an improved image using:
Method | CLIP | LPIPS* | VENDI | DINO | User Study-Q1 | User Study-Q2 |
LEDITS++ | 0.396 0.028 | 0.252 0.074 | 2.88 1.11 | 0.678 0.160 | 3.21 1.27 | 3.50 1.25 |
InstructPix2Pix | 0.379 0.032 | 0.289 0.126 | 1.94 0.61 | 0.704 0.189 | 3.14 1.23 | 3.59 1.17 |
SDEdit | 0.381 0.033 | 0.308 0.068 | 3.19 1.23 | 0.737 0.162 | 3.31 1.21 | 3.15 1.16 |
TurboEdit | 0.389 0.031 | 0.192 0.071 | 2.34 0.94 | 0.735 0.173 | 3.23 1.29 | 2.63 1.20 |
Ours (Editing) | 0.417 0.030 | 0.414 0.157 | 3.70 1.97 | 0.744 0.185 | 3.34 1.34 | 3.74 1.21 |
SDXL | 0.404 0.033 | 0.636 0.069 | 6.63 2.85 | N/A | 4.37 0.99 | 3.56 1.10 |
Flux | 0.359 0.048 | 0.650 0.088 | 5.84 2.69 | N/A | 4.11 1.27 | 3.00 1.24 |
ConceptLab | 0.334 0.055 | 0.663 0.076 | 10.38 2.27 | N/A | 3.40 1.52 | 3.18 1.30 |
Ours (Generation) | 0.360 0.052 | 0.709 0.057 | 10.44 2.15 | N/A | 4.32 0.99 | 4.16 1.01 |
The process iterates until , where is the maximum allowed iterations and is an intermediate creative image. Users can provide real-time instructions to enhance creativity, which the Prompt Architect integrates into the evolving prompt, while the Refinement Strategist ensures artistic coherence. The final goal is:
where is the final optimized image, achieved through iterative refinement of the refined current image, .

### v2

# CREA: A Collaborative Multi-Agent Framework for Creative Image Editing and Generation

@@ Abstract
Creativity in AI imagery remains a fundamental challenge, requiring not only the generation of visually compelling content but also the capacity to add novel, expressive, and artistically rich transformations to images. Unlike conventional editing tasks that rely on direct prompt-based modifications, creative image editing requires an autonomous, iterative approach that balances originality, coherence, and artistic intent. To address this, we introduce CREA, a novel multi-agent collaborative framework that mimics the human creative process. Our framework leverages a team of specialized AI agents who dynamically collaborate to conceptualize, generate, critique, and enhance images. Through extensive qualitative and quantitative evaluations, we demonstrate that CREA significantly outperforms state-of-the-art methods in diversity, semantic alignment, and creative transformation. To the best of our knowledge, this is the first work to introduce the task of creative editing.
@@ 1 Introduction
Generative AI has significantly transformed the field of image generation, producing high-quality visuals with remarkable detail and realism. Advances in diffusion models [rombach2022high, esser2024scaling], GANs [karras2019style], and retrieval-augmented techniques [blattmann2022retrieval, venkatesh2024context] have enabled powerful capabilities in content synthesis, making AI-driven image creation an essential tool for artists, designers, and various creative industries. These models have been widely applied in diverse tasks, including image-to-image translation [isola2017image, parmar2023zero, tumanyan2023plug], inpainting [rombach2022high, lugmayr2022repaint, ju2024brushnet], style transfer [karras2019style, gal2022image, ruiz2023dreambooth, zhang2024towards], and content-aware editing [meng2021sdedit, brooks2023instructpix2pix, brack2024ledits++, deutch2024turboedit], and opening up new frontiers in digital art, advertising, and entertainment, revolutionizing creative workflows.
@@
To address this limitation, we propose the task of creative image editing, where the goal is to modify images in a way that enhances their creative and artistic qualities with minimal user intervention. Unlike conventional editing tasks that focus on making explicit, text-driven modifications, creative editing aims to transform images into novel, aesthetically rich compositions in a disentangled way. Our approach mimics the complex process of creative image generation and editing as a collaborative team effort, drawing inspiration from human workflows where specialists iteratively refine ideas to achieve a shared artistic vision. To achieve this, we introduce a novel collaborative multi-agent framework, where agents, each with a distinct role, such as a Creative Director or Art Critic, work in synergy to conceptualize, generate and refine creative outputs by grounding them in well-established guidelines distilled from established principles in creativity literature [boden2004creative, guilford1950creativity, amabile1983social, ramachandran1999science]. This structured yet flexible approach enables the disentangled creation of highly diverse and imaginative images, ensuring both novelty and coherence at every stage of the generation process. Since creativity is a complicated task to assess and evaluate, our agents leverage the extensive knowledge in multimodal LLMs [achiam2023gpt] based on state-of-the-art creativity principles inspired by previous work [boden2004creative, guilford1950creativity, ramachandran1999science, amabile1983social, berlyne1973aesthetics]. Through both qualitative and quantitative evaluations, we demonstrate that our method consistently produces edits perceived as more creative and aesthetically pleasing compared to baseline methods. Our contributions are as follows:
@@
We demonstrate the versatility of our method across image editing and generation, as well as its potential for enhancing personalization workflows and creative video generation.
@@ 2 Related Work
Text-to-Image Models. Recent advances in diffusion models [ho2020denoising, rombach2022high, saharia2022photorealistic] have revolutionized text-to-image synthesis, enabling high-fidelity image generation guided by textual prompts. Models such as DALLE-3 [dalle3_2023], SDXL [podell2023sdxl], and Flux [flux2024] demonstrate the ability to generate visually compelling images based on textual prompts [witteveen2022investigating, wang2022diffusiondb, liu2022design]. While these models produce high-quality outputs, they lack a structured mechanism to enforce creativity principles in generation. Personalized generation approaches such as DreamBooth [ruiz2023dreambooth] and Textual Inversion [gal2022image] focus on fine-tuning text-to-image models for specific subjects or styles. However, these methods optimize for style and subject consistency rather than generating creative images.
@@
Creative Image Generation and Editing. Research on creative generation in AI has advanced through GANs [goodfellow2014generativeadversarialnetworks, karras2019style] and diffusion models, leveraging contrastive loss and diversity-based objectives to encourage novel synthesis beyond category constraints [elgammal2017can, heyrani2021creativegan, sbai2018design]. Recent works such as ProCreate [lu2024procreate] use energy-based repulsion to steer diffusion models away from reference images, while Inspiration Tree [vinker2023concept] employs hierarchical decomposition for conceptual hybridization. ConceptLab [richardson2024conceptlab] tackles creative concept synthesis through diffusion priors, iteratively enforcing constraints to generate novel category members. C3 [han2025enhancing] proposes a training-free approach for enhancing creativity by amplifying low-frequency feature maps in early layers of Stable Diffusion models. Despite their advancements, these methods require expensive retraining or optimization, and fail to generalize to broader concepts. For image editing, ControlNet [zhang2023adding] extends Stable Diffusion, introducing external conditioning signals for localized modifications while maintaining structural integrity. SDEdit [meng2021sdedit] and Blended Latent Diffusion [avrahami2023blended] refine diffusion-based editing for finer control, while InstructPix2Pix [brooks2023instructpix2pix] allows text-based transformations via user prompts. However, these approaches focus on structural fidelity rather than creativity-driven, minimal-guidance transformations.
@@
Large Language Models Large Language Models (LLMs) such as GPT-4 [achiam2023gpt] and PaLM [chowdhery2023palm] have demonstrated remarkable capabilities in natural language understanding and generative tasks. These models leverage transformer architectures [vaswani2017attention], which allows them to process large amounts of text and generate coherent and contextually relevant outputs. In creative domains, LLMs have been increasingly integrated into AI-driven artistic workflows, assisting with idea generation, structured prompt synthesis, and artistic guidance [mazzone2019art]. GenArtist [wang2024genartist] uses a multimodal single agent for general image generation and editing but does not address creative image editing and fails to leverage the collaborative nature of multiple agents for complex use cases. Several studies have explored LLMs in the context of creativity [lu2024llm, zhao2025assessing, li2024map, delorenzo2024creativeval, lidata, franceschelli2025creativity]; however, to the best of our knowledge, creative editing tasks have not been explicitly tackled before. Additionally, agentic frameworks, where AI systems exhibit autonomy in decision-making, iterative refinement, and adaptive goal-setting have not yet been systematically explored in the context of creative image generation or editing.
@@ 3 Methodology
Creativity is a multifaceted process that spans ideation, critique, refinement, and the delicate balancing of novelty with coherence. Traditional prompt‑to‑image pipelines handle this spectrum in a single static step, leaving users to face the burden of crafting elaborate prompts, testing variations, and manually disentangling edits. Such manual workflows are both expert‑intensive; they assume deep domain knowledge to articulate abstract creative facets, and time‑consuming, often requiring many trial‑and‑error iterations before the desired concept emerges. Our agentic framework replaces this bottleneck with a dynamic, collaborative loop that coordinates specialized AI agents such as a creative director, art critic, and refinement strategist to reason jointly and iteratively evolve an image. By structuring creativity as a multi‑agent dialogue, the system mirrors human studio practice: each role contributes focused expertise while shared memory maintains global coherence. The result is higher‑quality, more interpretable edits achieved with less human effort and lower computational overhead than a manual, prompt‑tweaking regimen. Fig. 2 offers an overview, and the sections that follow detail each component of the framework.
@@ 3.1 Multi-Agent Framework Design
Prompt Architect . Translates conceptual ideas into contrastive prompts [chia2023contrastive] for each creativity principle, merging them into a high-creativity prompt for the Generative Executor agent for image editing/generation and refining them iteratively based on agent feedback.
Generative Executor . This agent uses T2I models such as Flux and ControlNet for image generation or editing, dynamically selecting the appropriate diffusion model and parameters to ensure creative outputs.
Art Critic . The Art Critic evaluates the generated image based on the creativity principles, assigning a score for each criterion. Given that LLMs can approximate human judgment in subjective evaluations [zheng2023judging], this agent uses a multimodal LLM judge to ensure sensible evaluation.
@@
Each agent maintains a private memory and utilizes role-specific tools to share task status and relevant information, ensuring informed decision-making at every step. We acknowledge that an agentic pipeline introduces extra orchestration effort compared with a naive single-pass T2I call. However, an agentic design is critical for several aspects (i) modularity - each agent encapsulates a narrow skill (prompt refinement, consistency auditing, local editing) that can be swapped or upgraded independently; (ii) interpretability - inter-agent messages expose an explicit reasoning trace; and (iii) extensibility - new capabilities (e.g., style transfer, safety filtering) may be added to expand the scope of the framework. See Supplementary Material for detailed agent compositions.
@@ 3.2 Collaborative Agentic Synthesis
Creativity Principles. Our framework leverages six creativity principles, grounded in state-of-the-art creativity theories, to systematically evaluate and measure creative output. Originality, measuring novelty and uniqueness, is inspired by Boden’s Theory of Creativity [boden2004creative] and Guilford’s Divergent Thinking [guilford1950creativity] framework. Expressiveness, which captures emotional impact, is influenced by Amabile’s Model of Creativity [amabile1983social] and Ramachandran & Hirstein’s Laws of Aesthetics [ramachandran1999science]. Aesthetic Appeal, assessing composition and harmony, is grounded in Martindale’s Aesthetic Model [martindale1984pleasures] and Berlyne’s Aesthetic Theory [berlyne1973aesthetics]. Technical Execution, evaluating craftsmanship and skill, draws from Amabile’s Model and AI Creativity Frameworks [amabile1983social]. Unexpected Associations, reflecting surprise and ingenuity, is supported by the Geneplore Model [finke1996creative] and Boden’s Combinational Creativity [boden2004creative]. Finally, Interpretability & Depth, which considers exploration potential, is informed by Ramachandran’s Laws [ramachandran1999science] and the Geneplore Model [finke1996creative] (see Appendix for more details). These six principles are used as a creativity template for agents to assess and refine the outputs.
3.3 Pre-Generation Planning and Image Synthesis
@@
The pre-generation planning phase serves as a structured ideation stage where agents collaboratively establish a creative blueprint, before image generation begins. This phase involves three key agents: the Creative Director , Prompt Architect , and Generative Executor , who collaboratively devise the creativity prompt that is both creatively rich and technically viable. First, Creative Director, interprets the initial image , either user-provided or generated using the user-provided concept to formulate a creativity blueprint , capturing the core theme, stylistic interpretation, visual structure, and necessary constraints to balance artistic flexibility with semantic coherence (see Appendix for more details). Based on , the Prompt Architect synthesizes a set of contrastive prompts , each conditioned on a distinct creativity principle and are merged into a high-creativity fused prompt through Chain-of-Thought [wei2022chain] reasoning where CoT-Fusion extracts salient conceptual and stylistic attributes from each prompt and synthesizes them into a coherent, balanced, and conceptually rich formulation. The Generative Executor evaluates the feasibility of and determines T2I model-specific constraints-such as ControlNet conditioning scale, image guidance scale to anticipate the nuances of the generated blueprint. Once all agents reach a consensus, the prompt moves to the next phase for creative image generation.
@@
Image Generation: The Generative Executor, plays a dual role, either generating a creative image using a T2I model or performing disentangled creative editing on an existing image, whether initially generated () or user-provided (). The high-creativity structured prompt formulated during the pre-generation planning phase is then taken by the Generative Executor to synthesize the initial creative image, using a text-to-image diffusion model, as where represents model-specific parameters-such as guidance scale and ControlNet conditioning scale. The generated image, serves as the starting point for the creative editing and iterative refinement process as described in the following sections.
@@
Image Editing: If the user provides an image instead of generating one, or if an enhancement to is required, the Generative Executor performs disentangled creative edits using where is the ControlNet model used to perform disentangled edits, is the high-creativity editing prompt generated in the pre-generation planning phase and represents parameters of . After generation or editing, the resulting image progresses to the post-generation evaluation phase to investigate if further iterative refinements are necessary.
3.4 Post-Generation Evaluation
@@
The post-generation evaluation assesses the edited creative image, (or for generation) against the creativity template, to maximize creativity. This Critic and the Creative Director collaborate, with the Critic systematically evaluating the initial edited image, based on the creativity template, . The Critic utilizes the LLM-as-a-Judge to evaluate the edited image by assigning a creativity score, for each of the six creativity criteria on a 1-5 scale. The overall Creativity Index, is then computed as: .
@@
If the total creativity score, , where is a predefined threshold, the edited image is considered suboptimal in creativity. The Creative Director then reviews the Critic’s evaluation and may challenge its assessment if the assigned scores misalign with the creative blueprint, . Once a consensus is reached, is either finalized or progresses to further enhancement based on the final creativity score .
@@ 3.4.1 Self-Enhancement with Optional User-Guidance
While one round of editing produces optimal creative outputs, further refinement can enhance creativity. In the self-enhancement phase, all agents iteratively refine the edited image, with optional human intervention. Given a maximum of iterations, each edited image is evaluated by the Critic to compute its . When , the Refinement Strategist, identifies the low scoring creative dimensions and proposes a refinement plan to enhance the corresponding weak dimensions, which the Prompt Architect uses to refine the editing prompt, as follows where represents the prompt adjustment that amplify specific creativity dimensions based on their evaluation scores and is the refined creativity prompt. The refined prompt is used by the Generative Executor to regenerate an improved image using .
| Method | CLIP | LPIPS* | VENDI | DINO | FID | KID | User Study-Q1 | User Study-Q2 |
| LEDITS++ | 0.396 0.028 | 0.252 0.074 | 2.88 1.11 | 0.678 0.160 | 312.50 | 21.45 | 3.21 1.27 | 3.50 1.25 |
| InstructPix2Pix | 0.379 0.032 | 0.289 0.126 | 1.94 0.61 | 0.704 0.189 | 314.42 | 22.51 | 3.14 1.23 | 3.59 1.17 |
| SDEdit | 0.381 0.033 | 0.308 0.068 | 3.19 1.23 | 0.737 0.162 | 304.37 | 18.78 | 3.31 1.21 | 3.15 1.16 |
| TurboEdit | 0.389 0.031 | 0.192 0.071 | 2.34 0.94 | 0.735 0.173 | 320.53 | 24.79 | 3.23 1.29 | 2.63 1.20 |
| Ours (Editing) | 0.417 0.030 | 0.414 0.157 | 3.70 1.97 | 0.744 0.185 | 294.19 | 14.02 | 3.34 1.34 | 3.74 1.21 |
| SDXL | 0.404 0.033 | 0.636 0.069 | 6.63 2.85 | N/A | 282.22 | 9.67 | 4.37 0.99 | 3.56 1.10 |
| Flux | 0.359 0.048 | 0.650 0.088 | 5.84 2.69 | N/A | 270.69 | 9.94 | 4.11 1.27 | 3.00 1.24 |
| ConceptLab | 0.334 0.055 | 0.663 0.076 | 10.38 2.27 | N/A | 272.97 | 6.49 | 3.40 1.52 | 3.18 1.30 |
| Ours (Generation) | 0.360 0.052 | 0.709 0.057 | 10.44 2.15 | N/A | 248.67 | 5.94 | 4.32 0.99 | 4.16 1.01 |
The process iterates until , where is the maximum allowed iterations and is an intermediate creative image. Users can provide real-time instructions to enhance creativity, which the Prompt Architect integrates into the evolving prompt, while the Refinement Strategist ensures artistic coherence. The final goal is where is the final optimized image, achieved through iterative refinement of the refined current image, .